# Bogomolov property and Galois representations

### v1

@@ 1. Introduction
$=\left(\sum_{w\mid\infty}\frac{d_{n,w}}{d_{n}}\right)\log 2-\left(\sum_{w\in S}\frac{d_{n,w}}{d_{n}}\right)|H^{\prime}_{n}|\log p+2p^{\lambda}h(\beta^{\prime}).$
Since $\sum_{w\mid\infty}\frac{d_{n,w}}{d_{n}}=1$ and since $d_{n,w}=d_{n,v}$ for any place $w\mid p$ of $L_{n}$, and thus in particular for any $w\in S$, we get
(2.9) $2p^{\lambda}h(\beta^{\prime})\geq\frac{d_{n,v}}{d_{n}}|S|\cdot|H^{\prime}_{n}|\log p-\log 2$
To conclude, we need a lower bound for $S$. Note that $\sigma\in H^{\prime}_{n}\lhd\mathrm{Gal}(F_{0}L_{n}/{\mathbb{Q}})$, thus the orbit ${\mathcal{O}}$ of $\sigma$ is contained in $H^{\prime}_{n}$. Thus:
$|C(\sigma)|=\frac{d_{n}}{|{\mathcal{O}}|}\geq\frac{d_{n}}{|H^{\prime}_{n}|}.$
The stabiliser of $v$ under the action of $\mathrm{Gal}(F_{0}L_{n}/{\mathbb{Q}})$ is by definition the decomposition group $D(v|p)$ of $v$ over $p$ which has (again by definition) cardinality $d_{v}$. Thus,
$|S|=\frac{|C(\sigma)|}{|D(v|p)\cap C(\sigma)|}\geq\frac{d_{n}}{d_{n,v}}|H^{\prime}_{n}|^{-1}.$
@@
$c=\frac{\log(p/2)}{2p^{\lambda}}>0.$
As a toy example we deduce from Proposition 2.6 the following statement which generalize the main result of [2].
@@
Let $F/{\mathbb{Q}}$ be a Galois extension unramified at $p$ with property (B). Then $F({\boldsymbol{\mu}}_{p^{\infty}})$ has property (B) as well.
@@
Let $L={\mathbb{Q}}({\boldsymbol{\mu}}_{p^{\infty}})$. Since $L/{\mathbb{Q}}$ is totally ramified at $p$ and abelian, Assumption 2.2 is trivially satisfied. We consider the filtration given by $L_{n}={\mathbb{Q}}(\zeta_{p^{n}})$. Then it is easy to see that $G^{n}_{i}=\mathrm{Gal}({\mathbb{Q}}(\zeta_{p^{n}})/{\mathbb{Q}}(\zeta_{p^{j}}))$, for $p^{j-1}\leq i\leq p^{j}-1$ (see [19, Ch. II, §10, Ex. 1]). Thus, in the notation of Assumption (2.3), $e_{n}=(p-1)p^{n-1}$, $i_{n}=p^{n-1}-1$ and $H_{n}=G_{i_{n}}$. We then have $e_{n}/(i_{n}+1)\leq p$ and $|H_{n}|\leq p$. Clearly, $L_{n}^{H_{n}}=L_{n-1}$. The Assumption 2.3 therefore holds with $C_{1}=C_{2}=p$. ∎
3. Galois representation attached to a modular form.
For a field $F$, $\overline{F}$ denotes an algebraic closure of $F$, and $G_{F}=\mathrm{Gal}(\overline{F}/F)$ the absolute Galois group of $F$, endowed with the profinite topology.
Let $\Omega$ be a group and $\rho:G_{F}\to\Omega$ is a homomorphism; we shall denote by $F(\rho)=\overline{F}^{\ker\rho}$ the field cut out by $\rho$. Then $F(\rho)/F$ is a Galois extension whose Galois group is isomorphic to ${\rm Im}(\rho)$. Moreover if $\ker{\rho}$ is closed, then $G_{F(\rho)}\simeq\ker\rho$. Note that $\rho$ induces an isomorphism $\mathrm{Gal}(F(\rho)/F)\rightarrow{\rm Im}(\rho)$ which as usual we denote by the same letter $\rho$. We shall say that $\rho$ is unramified (resp. totally ramified) if $F(\rho)$ is.
From now on we fix a normalized new eigenform $f=\sum_{n}a_{n}q^{n}\in S_{k}(\Gamma_{0}(N))$ for the Hecke algebra of weight $k\geq 2$. We drop the index $f$ when it is clear from the context. We recall the relevant definitions. We denote by $K$ the number field generated by the $a_{n}$’s and by $\mathcal{O}$ its integer ring. A construction due to Shimura and Deligne associates to $f$ a family of compatible Galois representations:
$\rho_{v}:G_{\mathbb{Q}}\longrightarrow GL_{2}(\mathcal{O}_{v}),\quad v\hbox{ finite place of }K.$
[…]
(3.2) The characteristic polynomial of the image by ρv ofa Frobenius element at  with Npv is $\hbox{The characteristic polynomial of the image by $\rho_{v}$ of}\\ \hbox{a Frobenius element at $\ell$ with $\nmid Np_{v}$ is $X^{2}-a_{\ell}X+\ell^{k-1}$.}$
Remark 3.1.
Let $p$ be a prime and $v$ be any place of $K$ dividing $p$. Then, by (3.2) and by Chebotarev lemma, the determinant of $\rho_{v}$ is ${\varepsilon}_{p}^{k-1}$, where ${\varepsilon}_{p}\colon G_{\mathbb{Q}}\rightarrow{\mathbb{Z}}_{p}^{\times}$ is the $p$-adic cyclotomic character. (Indeed, for a Frobenius element Frobsubscriptnormal-$\mathrm{Frob}_{\ell}$ with $\ell\nmid Np$, we have: $\det\big{(}\rho_{v}(\mathrm{Frob}_{\ell})\big{)}=\ell^{k-1}={\varepsilon}_{p}^{k-1}(\mathrm{Frob}_{\ell})$.)
We let
$\rho=\prod_{v}\rho_{v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\hat{\mathcal{O}}),$
where $\hat{\mathcal{O}}$ is the profinite completion of $\mathcal{O}$. For a prime $p$ and $n>1$, we also denote by $\rho(p^{n})$ the reduction $G_{\mathbb{Q}}\to\mathrm{GL}_{2}(\hat{\mathcal{O}}/p^{n}\hat{\mathcal{O}})=\mathrm{GL}_{2}(\mathcal{O}/p^{n}\mathcal{O}$ of $\rho$.
Remark 3.2.
Let $p=\prod\limits_{v\mid p}{\mathfrak{P}}_{v}^{e_{v}}$ be a prime which does not divide $N$. Let $n\geq 1$ be an integer. Then
                 {tikzpicture}
                 {tikzpicture}
We let
We consider the following assumption on $p$:
Assumption 3.3.
  • $(P0)$
    $p\nmid N$
  • $(P1)$
    $a_{p}=0$
  • $(P2)$
    $\rho(p)$ has image $\widehat{G}(\mathcal{O}/p\mathcal{O})$
$\rho_{p}=\prod_{v|p}\rho_{v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\hat{\mathcal{O}})$
and
$\rho^{(p)}=\prod_{v\nmid p}\rho_{v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\hat{\mathcal{O}}).$
Theorem 3.4.
If Assumption 3.3 (P0) and (P1) hold, then $\rho^{(p)}$ has property (B).
Proof.
Let $F=\mathbb{Q}(\rho^{(p)})$. By our assumptions and by (3.1) and (3.2), $F$ is unramified at $p$ and the characteristic polynomial of $\rho^{(p)}(\mathrm{Frob}_{p})$ is $𝑘1X^{2}+p^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{k-1}}$. It follows that $\mathrm{Frob}_{p}^{2}$ lies in the center of $\mathrm{Gal}(F/{\mathbb{Q}})$. Then Property (B) for $F$ is ensured by Proposition 2.1. ∎
Corollary 3.5.
Theorem 1.3 implies Theorem 1.4.
Proof.
Let again $F=\mathbb{Q}(\rho^{(p)})$. then $F$ is unramified at $p$, and it has Property (B) by Theorem 3.4. If Theorem 1.3 is true, then $F(\rho_{p})$ has Property (B). But
$F(\rho)=\mathbb{Q}(\rho^{(p)}\times\rho_{p})=\mathbb{Q}(\rho).$
4. Crystalline representation.
@@
In what follows, we recall some results on crystalline representations. We fix a prime $p$ and an embedding ¯¯$\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}}_{p}$, which allows to identify $G_{{\mathbb{Q}}_{p}}=\mathrm{Gal}(\overline{\mathbb{Q}}_{p}/{\mathbb{Q}}_{p})$ with a subgroup of $G_{\mathbb{Q}}$. Note that this is helpful but harmful in our context, as some results may depend on the chosen embedding. To simplify notations and to avoid future conflicts, we simply denote by ${\varepsilon}$ the $p$-adic cyclotomic character ${\varepsilon}_{p}$. Let $q=p^{2}$ and denote by ${\mathbb{Q}}_{q}$ the unique unramified quadratic extension of ${\mathbb{Q}}_{p}$. Let $v$ be a place of $K$ over $p$. The two dimensional irreducible crystalline representations of $G_{\mathbb{Q}_{p}}$ have been classified in [6, Proposition 3.1]. They are of the form $V_{k^{\prime},a,\chi}=V_{k^{\prime},a}\otimes\chi$, where $k^{\prime}$ is an integer $\geq 2$, $a$ is an element in the maximal ideal of the valuation ring of ¯$\overline{\mathbb{Q}}_{p}$, and $\chi$ is a crystalline character, that is the product of an unramified character by an integral power of the cyclotomic character. The only isomorphisms between them are $V_{k^{\prime},a,\chi}\simeq V_{k^{\prime},-a,\chi\mu_{-1}}$, where $\mu_{-1}$ is the unramified quadratic character of $G_{\mathbb{Q}_{p}}$. Up to a twist by a power of the cyclotomic character, we can suppose that $\chi$ is unramified. In this case $V_{k^{\prime},a,\chi}$ has Hodge-Tate weights $(0,k^{\prime}-1)$, and the $\varphi$-filtered module $D=D_{k^{\prime},a,\chi}$ associated to $V_{k^{\prime},a,\chi}$ is ¯$\overline{\mathbb{Q}}_{p}e_{1}+\overline{\mathbb{Q}}_{p}e_{2}$ with $\varphi$-action given by
@@
In particular, the characteristic polynomial of $\varphi$ on the associated $\varphi$-filtered module is equal to $X^{2}+aX+\chi(p^{-1})^{2}p^{k^{\prime}-1}$.
Let’s come back to our representations $\rho_{v}$, with $v\mid p$. Let as assume $p\nmid N$ (Assumption 3.3 (P0)). By [23, Theorem 1.2.4 (ii)], ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is crystalline and the characteristic polynomial of $\varphi$ on the associated $\varphi$-filtered module is equal to $X^{2}-a_{p}X+p^{k-1}$. Moreover, representations associate to modular forms of weight $k$ have Hodge-Tate weight $(0,k-1)$, (see [11]). It follows that the filtered module associated to ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is $D_{k^{\prime},a,\chi}$ with $k^{\prime}=k$, $a=a_{p}$, and $\chi(p)=1$ or $-1$. Now assume also $a_{p}=0$ (Assumption 3.3 (P1)). Since $V_{k^{\prime},a,\chi}\simeq V_{k^{\prime},-a,\chi\mu_{-1}}$, where $\mu_{-1}$ is the unramified quadratic character of $G_{{\mathbb{Q}}_{p}}$, in both cases $\rho_{v}|_{G_{\mathbb{Q}_{p}}}$ is isomorphic to $V_{k,0}$. Taking into account the explicit description of $V_{k,0}$ in [6, Prop. 3.2] we obtain:
@@
Let us assume 3.3 (P0) and (P1). Then for every prime $v$ of $\mathcal{O}$ above $p$, ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is isomorphic to the crystalline representation
@@
We shall also make use of the following proposition.
@@
Let us assume 3.3 (P0)-(P1) and $k\leq p+1$. Let $v$ be a place of $K$ over $p$. Then, up to conjugation by an element of $\mathrm{GL}_{2}(\mathcal{O}_{v})$, the restriction ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ takes values in $\mathrm{GL}_{2}({\mathbb{Z}}_{p})$.
@@
Let $\mathfrak{P}$ be the prime ideal corresponding to $v$. Denote by $\rho_{v}(\mathfrak{P})$ the reduction $G_{\mathbb{Q}}\to\mathrm{GL}_{2}(\mathcal{O}_{v}/\mathfrak{P}\mathcal{O}_{v})=\mathrm{GL}_{2}(\mathcal{O}/\mathfrak{P})$ of $\rho_{v}$. Since $12\leq k\leq p+$ 1, by an unpublished result by Fontaine (see [10, Theorem 2.6]), the restriction of $\overline{\rho_{v}(\mathfrak{P})}$ to a decomposition group [^2] By Proposition 4.1 this restriction does not depend on the decomposition group but only on $p$. at $p$ is absolutely irreducible, so that $\rho_{v}$ a fortiori is absolutely irreducible. By Proposition 4.1, the representations ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}\simeq V(k,0)$ take values in $\mathrm{GL}_{2}(\mathbb{Q}_{p})$, so that the traces lie in $\mathbb{Z}_{p}$. By [7, Theorem 1], up to conjugation by an element of $\mathrm{GL}_{2}(\mathcal{O}_{v})$, the restriction ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ takes values in $\mathrm{GL}_{2}(\mathbb{Z}_{p})$. ∎
@@
Let $p$ be a rational prime satisfying Assumptions 3.3 (P0), (P1). By Propositions 4.1, for every prime $v$ of $\mathcal{O}$ above $p$, ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is a twist by an unramified character of the induced representation of a power of a character arising from the Tate module of a Lubin-Tate formal group over $\mathbb{Z}_{q}$ (see [9, Théorème I.2.1]). This is exactly the same situation occuring for the representation of $G_{{\mathbb{Q}}_{p}}$ on the $p$-adic Tate module of an elliptic curve $E$ over ${\mathbb{Q}}$ having $p$ as a supersingular prime, and this is the key point of the fundamental Galois theoretic analysis of the $p$-power torsion points on $E$ in [15, Lemma 3.3]. The proof of this lemma rests on the theory of Lubin-Tate modules and its relation to local class field theory and thus only depends on the local $p$-adic representation. We provide here a reformulation suitable for our situation, without any reference to the theory of elliptic curves.
@@
If $\nu$ surjective ans the extensions $L(\nu)/L$, $L(\omega)/L$ are linearly disjoint, then the product $\nu\omega$ is surjective and the restriction
@@
This formula implies that $\nu$ is surjective when restricted to $\mathrm{Gal}(\overline{L}/L(\omega))=\ker\omega$. Since $\nu=\nu\omega$ over $\ker\omega$, this prove that $\nu\omega$ is surjective. We now prove the second assertion. The restriction (4.1) is clearly injective. Let $\tau\in\mathrm{Gal}(L(\omega)/L))$. Since $\nu$ is surjective, there exists an element $\sigma\in\mathrm{Gal}(L(\nu)L(\omega)/L)$ such that $\nu(\sigma)=\omega(\tau)^{-1}$ and $\sigma$ restricts to $\tau$ in $\mathrm{Gal}(L(\omega)/L)$. Then $\nu(\sigma)\omega(\sigma)=1$ so that $\sigma\in\mathrm{Gal}(L(\nu)L(\omega)/L(\nu\omega))$. ∎
@@
If $\nu$ is surjective and totally ramified and $\omega$ unramified with finite image, then
@@
[…]
@@
Let us prove (i) and (ii). Since $\nu$ is totally ramified and $\omega$ is unramified, the two extensions $L(\nu)/L$ and $L(\omega)/L$ are linearly disjoint. Lemma 4.4 then shows that $\nu\omega$ is surjective and
@@
From now on, we shall make the following
Assumption 4.6.
k=2.
The following is an analogue of [15, Lemma 3.3].
@@
    The extension ${\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q}$ is totally ramified and abelian of degree $q^{n-1}$. Moreover
    $\mathrm{Gal}\left({\mathbb{Q}}_{q}(p^{n})/{\mathbb{Q}}_{q}\right)\simeq{\mathbb{Z}}/(q-1){\mathbb{Z}}\times\left({\mathbb{Z}}/p^{n-1}{\mathbb{Z}}\right)^{2}$

### v3

@@ 2. Notations and preliminary results
$=\left(\sum_{w\mid\infty}\frac{d_{n,w}}{d_{n}}\right)\log 2-\left(\sum_{w\in S}\frac{d_{n,w}}{d_{n}}\right)|\overline{H^{\prime}_{n}}|\log p+2p^{\lambda}h(\beta^{\prime}).$
Since $\sum_{w\mid\infty}\frac{d_{n,w}}{d_{n}}=1$ and since $d_{n,w}=d_{n,v}$ for any place $w\mid p$ of ${\mathbb{L}}_{n}$, and thus in particular for any $w\in S$, we get
(3.10) $2p^{\lambda}h(\beta^{\prime})\geq\frac{d_{n,v}}{d_{n}}|S|\cdot|\overline{H^{\prime}_{n}}|\log p-\log 2$
To conclude, we need a lower bound for $S$. Note that $\sigma\in\overline{H^{\prime}_{n}}\lhd\mathrm{Gal}({\mathbb{F}}_{0}{\mathbb{L}}_{n}/{\mathbb{Q}})$, thus the orbit $O$ of $\sigma$ is contained in $\overline{H^{\prime}_{n}}$. Thus:
$|C(\sigma)|=\frac{d_{n}}{|O|}\geq\frac{d_{n}}{|\overline{H^{\prime}_{n}}|}.$
The stabiliser of $v$ under the action of $\mathrm{Gal}({\mathbb{F}}_{0}{\mathbb{L}}_{n}/{\mathbb{Q}})$ is by definition the decomposition group $D(v|p)$ of $v$ over $p$ which has (again by definition) cardinality $d_{v}$. Thus,
$|S|=\frac{|C(\sigma)|}{|D(v|p)\cap C(\sigma)|}\geq\frac{d_{n}}{d_{n,v}}|\overline{H^{\prime}_{n}}|^{-1}.$
@@
$c=\frac{\log(p/2)}{2p^{\lambda}}>0.$ $\hfill\qed$
As a toy example we deduce from Proposition 3.4 the following statement which generalizes the main result of [2].
@@
Let ${\mathbb{F}}/{\mathbb{Q}}$ be a Galois extension unramified at $p$ with property (B). Then ${\mathbb{F}}({\boldsymbol{\mu}}_{p^{\infty}})$ has property (B) as well.
@@
Let ${\mathbb{L}}={\mathbb{Q}}({\boldsymbol{\mu}}_{p^{\infty}})$. Since ${\mathbb{L}}/{\mathbb{Q}}$ is totally ramified at $p$ and abelian, Assumption 3.2 is trivially satisfied. We consider the filtration given by ${\mathbb{L}}_{n}={\mathbb{Q}}(\zeta_{p^{n}})$. Then it is easy to see that $G^{n}_{i}=\mathrm{Gal}({\mathbb{Q}}(\zeta_{p^{n}})/{\mathbb{Q}}(\zeta_{p^{j}}))$, for $p^{j-1}\leq i\leq p^{j}-1$ (see [19, Ch. II, §10, Ex. 1]). Thus, in the notation of Assumption (3.3), $e_{n}=(p-1)p^{n-1}$, $i_{n}=p^{n-1}-1$ and $\overline{H_{n}}=G_{i_{n}}^{n}$. We then have $e_{n}/(i_{n}+1)\leq p$ and $|\overline{H_{n}}|\leq p$. Clearly, ${\mathbb{L}}_{n}^{\overline{H_{n}}}={\mathbb{L}}_{n-1}$. The Assumption 3.3 therefore holds with $C_{1}=C_{2}=p$. ∎
4. The local representation at $v\mid p$.
@@
In what follows, we recall some results on crystalline representations. We fix a prime $p$ and an embedding ¯¯$\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}}_{p}$, which allows to identify $G_{{\mathbb{Q}}_{p}}=\mathrm{Gal}(\overline{\mathbb{Q}}_{p}/{\mathbb{Q}}_{p})$ with a subgroup of $G_{\mathbb{Q}}$. Note that this is helpful but harmful in our context, as some results may depend on the chosen embedding. To simplify notations and to avoid future conflicts, we simply denote by ${\varepsilon}$ the $p$-adic cyclotomic character ${\varepsilon}_{p}$. Let $q=p^{2}$ and denote by ${\mathbb{Q}}_{q}$ the unique unramified quadratic extension of ${\mathbb{Q}}_{p}$. Let $v$ be a place of ${\mathbb{K}}$ over $p$. The two dimensional irreducible crystalline representations of $G_{\mathbb{Q}_{p}}$ have been classified in [6, Proposition 3.1]. They are of the form $V_{k^{\prime},a,\chi}=V_{k^{\prime},a}\otimes\chi$, where $k^{\prime}$ is an integer $\geq 2$, $a$ is an element in the maximal ideal of the valuation ring of ¯$\overline{\mathbb{Q}}_{p}$, and $\chi$ is a crystalline character, that is the product of an unramified character by an integral power of the cyclotomic character. The only isomorphisms between them are $V_{k^{\prime},a,\chi}\simeq V_{k^{\prime},-a,\chi\mu_{-1}}$, where $\mu_{-1}$ is the unramified quadratic character of $G_{\mathbb{Q}_{p}}$. Up to a twist by a power of the cyclotomic character, we can suppose that $\chi$ is unramified. In this case $V_{k^{\prime},a,\chi}$ has Hodge-Tate weights $(0,k^{\prime}-1)$, and the $\varphi$-filtered module $D=D_{k^{\prime},a,\chi}$ associated to $V_{k^{\prime},a,\chi}$ is ¯$\overline{\mathbb{Q}}_{p}e_{1}+\overline{\mathbb{Q}}_{p}e_{2}$ with $\varphi$-action given by
@@
In particular, the characteristic polynomial of $\varphi$ on the associated $\varphi$-filtered module is equal to $X^{2}-aX+\chi(p^{-1})^{2}p^{k^{\prime}-1}$.
Let’s come back to our representations $\rho_{v}$, with $v\mid p$. Let as assume $p\nmid N$ (Assumption 1.3 $(P0)$). By [25, Theorem 1.2.4 (ii)], ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is crystalline and the characteristic polynomial of $\varphi$ on the associated $\varphi$-filtered module is equal to $X^{2}-a_{p}X+p^{k-1}$. Moreover, representations associate to modular forms of weight $k$ have Hodge-Tate weight $(0,k-1)$, (see [11]). It follows that the filtered module associated to ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is $D_{k^{\prime},a,\chi}$ with $k^{\prime}=k$, $a=a_{p}$, and $\chi(p)=1$ or $-1$. Now assume also $a_{p}=0$ (Assumption 1.3 $(P1)$). Since $V_{k^{\prime},a,\chi}\simeq V_{k^{\prime},-a,\chi\mu_{-1}}$, where $\mu_{-1}$ is the unramified quadratic character of $G_{{\mathbb{Q}}_{p}}$, in both cases $\rho_{v}|_{G_{\mathbb{Q}_{p}}}$ is isomorphic to $V_{k,0}$. Taking into account the explicit description of $V_{k,0}$ in [6, Prop. 3.2] we obtain:
@@
Let us assume 1.3 $(P0)$ and $(P1)$. Then for every prime $v$ of $\mathcal{O}$ above $p$, ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is isomorphic to the crystalline representation
@@
We also need the following proposition.
@@
Let us assume 1.3 $(P0)$-$(P1)$-$(P3)$. Let $v$ be a place of ${\mathbb{K}}$ over $p$. Then, up to conjugation by an element of $\mathrm{GL}_{2}(\mathcal{O}_{v})$, the restriction ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ takes values in $\mathrm{GL}_{2}({\mathbb{Z}}_{p})$.
@@
Let $\mathfrak{P}$ be the prime ideal corresponding to $v$. Denote by $\rho_{v}(\mathfrak{P})$ the reduction $G_{\mathbb{Q}}\to\mathrm{GL}_{2}(\mathcal{O}_{v}/\mathfrak{P}\mathcal{O}_{v})=\mathrm{GL}_{2}(\mathcal{O}/\mathfrak{P})$ of $\rho_{v}$. Since $p+1\nmid k-1$ (which holds a fortiori since $\frac{p+1}{2}\nmid k-1$ by Assumption 1.3 $(P3)$), [7, Lemma 2.1 (1)]) ensures that the restriction of $\rho_{v}(\mathfrak{P})$ to a decomposition group [^4] By Proposition 4.1 this restriction does not depend on the decomposition group but only on $p$. at $p$ is absolutely irreducible, so that $\rho_{v}$ a fortiori is absolutely irreducible. By Proposition 4.1, the representations ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}\simeq V(k,0)$ take values in $\mathrm{GL}_{2}(\mathbb{Q}_{p})$, so that the traces lie in $\mathbb{Z}_{p}$. By [8, Theorem 1], up to conjugation by an element of $\mathrm{GL}_{2}(\mathcal{O}_{v})$, the restriction ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ takes values in $\mathrm{GL}_{2}(\mathbb{Z}_{p})$. ∎
@@
Let $p$ be a rational prime satisfying Assumptions 1.3 $(P0)$, $(P1)$. By Propositions 4.1, for every prime $v$ of $\mathcal{O}$ above $p$, ${\rho_{v}}_{|G_{{\mathbb{Q}}_{p}}}$ is a twist by an unramified character of the induced representation of a power of a character arising from the Tate module of a Lubin-Tate formal group over $\mathbb{Z}_{q}$ (see [10, Théorème I.2.1]). This is exactly the same situation occuring for the representation of $G_{{\mathbb{Q}}_{p}}$ on the $p$-adic Tate module of an elliptic curve ${\mathcal{E}}$ over ${\mathbb{Q}}$ having $p$ as a supersingular prime, and this is the key point of the fundamental Galois theoretic analysis of the $p$-power torsion points on $E$ in [15, Lemma 3.3]. The proof of this lemma rests on the theory of Lubin-Tate modules and its relation to local class field theory and thus only depends on the local $p$-adic representation. We provide here a reformulation suitable for our situation, without any reference to the theory of elliptic curves.
@@
Assume that $\mathrm{im}(\omega)\subseteq\mathrm{im}(\nu)$ and the extensions $L(\nu)/L$, $L(\omega)/L$ are linearly disjoint; then $\mathrm{im}(\nu\omega)=\mathrm{im}(\nu)$ and the restriction
@@
This formula implies that the image of $\nu$ restricted to $\mathrm{Gal}(\overline{L}/L(\omega))=\ker\omega$ coincides with $\mathrm{im}(\nu)$. Since $\nu=\nu\omega$ over $\ker\omega$, this prove that $\mathrm{im}(\nu\omega)=\mathrm{im}(\nu)$. We now prove the second assertion. The restriction (4.1) is clearly injective. Let $\tau\in\mathrm{Gal}(L(\omega)/L))$. Since $\mathrm{im}(\omega)\subseteq\mathrm{im}(\nu)$, there exists an element $\sigma\in\mathrm{Gal}(L(\nu)L(\omega)/L)$ such that $\nu(\sigma)=\omega(\tau)^{-1}$ and $\sigma$ restricts to $\tau$ in $\mathrm{Gal}(L(\omega)/L)$. Then $\nu(\sigma)\omega(\sigma)=1$ so that $\sigma\in\mathrm{Gal}(L(\nu)L(\omega)/L(\nu\omega))$. ∎
@@
Assume that $\mathrm{im}(\omega)\subseteq\mathrm{im}(\nu)$. If $\nu$ is totally ramified and $\omega$ unramified with finite image, then
@@
$L$$L(\nu\omega)$$L(\omega)$$L(\nu)$$E:=L(\nu\omega)L(\omega)=L(\nu)L(\omega)$ finite unr tot. ram. finite unr finite unr
@@
Let us prove $(i)$ and $(ii)$. Since $\nu$ is totally ramified and $\omega$ is unramified, the two extensions $L(\nu)/L$ and $L(\omega)/L$ are linearly disjoint. Lemma 4.4 then shows that ${\mathrm{im}}(\nu\omega)={\mathrm{im}}(\nu)$ and
@@
We are now ready to completely describe the ramification of the local representation $\rho_{p}|_{G_{{\mathbb{Q}}_{p}}}$, as in [15, Lemma 3.3]. Recall that $q=p^{2}$. We put:
$\delta=\frac{q-1}{(q-1,k-1)}.$
Notice that $\delta\not=1$ because $k$ is even.
@@
Let $n\geq 1$. We denote by $\epsilon_{2,n}$ the composition of $\epsilon_{2}$ with the reduction modulo $p^{n}$.
@@
    The extension ${\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q}$ is totally ramified and abelian of degree $\delta q^{n-1}$, and
    $\mathrm{Gal}({\mathbb{Q}}_{q}(\epsilon_{2,n}^{k-1})/{\mathbb{Q}}_{q})\simeq\{g^{k-1}\,|\,g\in({\mathbb{Z}}_{q}/p^{n}{\mathbb{Z}}_{q})^{\times}\}\simeq{\mathbb{Z}}/\delta{\mathbb{Z}}\times\left({\mathbb{Z}}/p^{n-1}{\mathbb{Z}}\right)^{2}.$